Grade 12Mechanics

3 rods each of mass m and lenght l are joined together to form equilateral triangle.find moment of inertia

Pranav Neelakandan
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
 
I=I+I+I 
Now
I=(IcM)+md² 
=ML²/12+M(L/2√3)² 
=ML²/12+ML²/4×3 
=ML²/12+ML²/12 
=2ML²/12 
=ML²/6 
=3×ML²/6 
=ML²/2
 
Regards
Arun (AskIITians Forum Expert)